OpenStudy (kt7779):

i need help i give metals!!!! ratio 4:6 = ? : 7.5 please explain :)

jimthompson5910 (jim_thompson5910):

4/6 = x/7.5 4*7.5 = 6x 30 = 6x x = ???

OpenStudy (kt7779):

i dont know what x =!

jimthompson5910 (jim_thompson5910):

well if 6x = 30 then x = ???

jimthompson5910 (jim_thompson5910):

put another way 6 times some number = 30 what is that "some number" ?

OpenStudy (kt7779):

5

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (kt7779):

thx soooo much!

jimthompson5910 (jim_thompson5910):

so the ratio is 4:6 = 5:7.5

OpenStudy (kt7779):

omg thx so much i an a fan now!

jimthompson5910 (jim_thompson5910):

you're welcome, glad to be of help

OpenStudy (kt7779):

thx again!

jimthompson5910 (jim_thompson5910):

sure thing